OpenStudy (anonymous):

The length of a rectangle is six times its width. If the area of the rectangle is 600 cm2 , find its perimeter.

OpenStudy (anonymous):

To make this problem easier, draw a rectangles. Label the length 6w because it is 6 times the width. Label width w. Then set up an equation that says length times width = area. This is the equation: 6w*w=384. Multiply, you get 6w^2=384, divide both sides by six, you get w^2=64, then take the square root of both sides, so sqrtw^2=sqrt64. x=8. Width is 8 and the length is 48! Then add all sides, 8+8+48+48=112.

OpenStudy (anonymous):

im a nerd

OpenStudy (anonymous):

since when is \(8\times 48=600\) ?

OpenStudy (anonymous):

copy and paste from some other problem i guess, but the idea is right equation should be \[6w^2=600\] making \(w^2=100\) and so \(w=10\)

OpenStudy (anonymous):

The perimeter is: 2 w + 2 (6 w) at w = 10, 20 + 120 = 140

OpenStudy (anonymous):

Thank you ! Can you please help me with another one?

OpenStudy (anonymous):

y you block me